OpenStudy (anonymous):

solve for x: x/3 + x/4 = 21/12

OpenStudy (anonymous):

The common denominator is 12. Multiply first expression on 4, second on 3. 4x + 3x / 12 = 21/12. Multiply both by 12. 7x = 21, x = 3